# A transversality theorem for semi-algebraic sets with application to signal recovery from the second moment and cryo-EM

### v2

@@ Appendix A A transversality theorem for complex representations
When $\Theta=\mathrm{GL}(V)$, then two vectors $x,y$ are equivalent if and only if $y=e^{\iota\theta}x$ and as in the proof of Theorem 2.2, $\Theta(x,y)$ is empty unless they have the same magnitude, in which case they are equivalent. Likewise, if $x,y$ are linearly independent, then we can take them to be the first two vectors in an ordered basis for $V$. If we represent $A\in\mathrm{GL}(V)$ as a matrix with respect to this ordered basis, the condition that $Ax=h\cdot Ay$ implies that the first row of $A$ lies in a real algebraic set of dimension at most $k(H)$.
@@
$S_{1}$ =$=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T=\lambda(y+T),|\lambda|=1\}$
$S_{2}$ =$=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T,(y+T)\text{ are linearly dependent but }|x+T|\neq|y+T|\}$
$S_{3}$ =$=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T\text{ and }y+T\text{ are linearly independent}\}.$
For each $\lambda\in S^{1}$ there is a unique translation $T$ such that $(x+T)=\lambda(y+T)$. Thus, we have that $\dim_{\mathbb{R}}S_{1}=\dim_{\mathbb{R}}(\mathrm{Aff}(V))-\dim_{\mathbb{R}}(V)+1\leq\dim_{\mathbb{R}}(\mathrm{Aff}(V)-K+1$. As in the real case $S_{2}$ is empty because the action of $H$ is norm preserving. Likewise, the same argument used in the real case shows that
$\dim S_{3}\leq\dim_{\mathbb{R}}\mathrm{Aff}(V)-(\dim_{\mathbb{R}}(V)-k(H))=\dim\mathrm{Aff}(V)-K,$
@@
Once again $U(x,y)$ is empty if $x,y$ are linearly dependent but not equivalent (i.e., they have different norms). If $x,y$ are linearly independent, then using the notation from the proof of Theorem 2.3, if $A\in U(x,y)$ then the second row $w_{2}$ of $A$ lies in the real algebraic subset $-ab^{-1}w_{1}+b^{-1}Hw_{1}$ for some constants $a,b$ with $a\neq 0$ and here $w_{1}$ is the first row. This means that given $w_{1}$, $w_{2}$ is taken from a real algebraic subset of $U(N)$ of dimension $k(H)$, which has real codimension at least $32N-3-k(H)=K-$ 3 in the set of unit vectors $w_{2}$ orthogonal to $w_{1}$. Once again we do not impose any further conditions on the rows $A$.
@@
imposes at least one additional real algebraic condition on $hw_{1}$ unless $a=0$ and $|b|=1$. Thus, if $x,y$ are not orthogonal vectors with the same magnitude, then the condition that $w_{2}$ is a unit vector shows that $w_{2}$ is taken from a real algebraic subset of $S^{2N-3}$ of dimension at most $k(H)-1$, so in this case $\dim U(x,y)\leq\dim U(V)-K+2$.

### v3

@@ Appendix A A transversality theorem for complex representations
When $\Theta=\mathrm{GL}(V)$, then two vectors $x,y$ are equivalent if and only if $y=e^{\iota\theta}x$ and as in the proof of Theorem 2.2, $\Theta(x,y)$ is empty unless they have the same magnitude, in which case they are equivalent. Likewise, if $x,y$ are linearly independent, then we can take them to be the first two vectors in an ordered basis for $V$. If we represent $A\in\mathrm{GL}(V)$ as a matrix with respect to this ordered basis, the condition that $Ax=h\cdot Ay$ implies that the first column of $A$ lies in a real algebraic set of dimension at most $k(H)$.
@@
$S_{1}$ $=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T=\lambda(y+T),|\lambda|=1\}$
$S_{2}$ $=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T,(y+T)\text{ are linearly dependent but }|x+T|\neq|y+T|\}$
$S_{3}$ $=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T\text{ and }y+T\text{ are linearly independent}\}.$
For each $\lambda\in S^{1}$ there is a unique translation $T$ such that $(x+T)=\lambda(y+T)$. Thus, we have that $\dim_{\mathbb{R}}S_{1}=\dim_{\mathbb{R}}\mathrm{Aff}(V)-\dim_{\mathbb{R}}(V)+1\leq\dim_{\mathbb{R}}\mathrm{Aff}(V)-K+1$. As in the real case $S_{2}$ is empty because the action of $H$ is norm preserving. Likewise, the same argument used in the real case shows that
$\dim S_{3}\leq\dim_{\mathbb{R}}\mathrm{Aff}(V)-(\dim_{\mathbb{R}}(V)-k(H))=\dim_{\mathbb{R}}\mathrm{Aff}(V)-K,$
@@
Once again $U(x,y)$ is empty if $x,y$ are linearly dependent but not equivalent (i.e., they have different norms). If $x,y$ are linearly independent, then using the notation from the proof of Theorem 2.3, if $A\in U(x,y)$ then the second column $w_{2}$ of $A$ lies in the real algebraic subset $ab^{-1}w_{1}+b^{-1}Hw_{1}$ for some constants $a,b$ with $a\neq 0$ and here $w_{1}$ is the first column. This means that given $w_{1}$, $w_{2}$ is taken from a real algebraic subset of $U(N)$ of dimension $k(H)$, which has real codimension at least 2N3k(H)=K32N-3-k(H)=K-3 in the set of unit vectors $w_{2}$ orthogonal to $w_{1}$. Once again we do not impose any further conditions on the columns $A$.
@@
imposes at least one additional real algebraic condition on $hw_{1}$ unless $a=0$ and $|b|=1$. Thus, if $x,y$ are not orthogonal vectors with the same magnitude, then the condition that $w_{2}$ is a unit vector shows that $w_{2}$ is taken from a real algebraic subset of $S^{2N-3}$ of dimension at most k(H)1k(H)-1, so in this case $\dim U(x,y)\leq\dim U(V)-K+2$.